find all second partial derivatives of

f(xy)=xcosy+ye^x


1
Expert's answer
2021-10-20T01:55:05-0400
"f(x,y)=x\cos y+ye^x"

"f_x=\cos y+ye^x"

"f_y=-x\sin y+e^x"



"f_{xx}=ye^x"

"f_{xy}=-\sin y+e^x=f_{yx}"

"f_{yy}=-x\cos y"